# Towards Generalised Pre-Training of Graph Models

### v3

@@ A.2 Exploration of Embedding Space
A.3 Loss Curves
Appendix B Further Results
B.1 Linear Transfer
@@ B.2 Transfer Learning
Table 8: Scores for fine-tuning models pre-trained on non-molecules, each with one non-molecular pre-training dataset excluded, with (Model-features) and without (Model) features. For the classification datasets we report AUROC and for regression datasets RMSE. Some datasets, where results are very similar for each model, are excluded. Bold text indicates a result performs equally (or better) than the model pre-trained on molecules (Chem).
Social (by excluded pre-training dataset)
Chem Twitch Neurons Cora Facebook Roads
molfreesolv 3.89 $\pm$ 0.21 4.14 $\pm$ 0.07 4.09 $\pm$ 0.02 4.17 $\pm$ 0.03 3.97 $\pm$ 0.11 4.24 $\pm$ 0.16
molesol 1.44 $\pm$ 0.09 1.37 $\pm$ 0.08 1.64 $\pm$ 0.06 1.67 $\pm$ 0.04 1.58 $\pm$ 0.13 1.59 $\pm$ 0.07
mollipo 1.04 $\pm$ 0.02 1.00 $\pm$ 0.01 0.99 $\pm$ 0.01 1.01 $\pm$ 0.01 1.00 $\pm$ 0.01 1.01 $\pm$ 0.01
molclintox 0.51 $\pm$ 0.03 0.48 $\pm$ 0.12 0.44 $\pm$ 0.03 0.43 $\pm$ 0.00 0.43 $\pm$ 0.01 0.54 $\pm$ 0.10
molbbbp 0.54 $\pm$ 0.04 0.56 $\pm$ 0.04 0.40 $\pm$ 0.02 0.60 $\pm$ 0.02 0.62 $\pm$ 0.01 0.51 $\pm$ 0.02
molbace 0.63 $\pm$ 0.13 0.73 $\pm$ 0.04 0.62 $\pm$ 0.03 0.71 $\pm$ 0.05 0.70 $\pm$ 0.04 0.69 $\pm$ 0.02
molhiv 0.54 $\pm$ 0.13 0.42 $\pm$ 0.06 0.65 $\pm$ 0.07 0.54 $\pm$ 0.15 0.66 $\pm$ 0.04 0.61 $\pm$ 0.06
Facebook 1.17 $\pm$ 0.17 0.29 $\pm$ 0.02 0.14 $\pm$ 0.00 0.84 $\pm$ 0.01 0.14 $\pm$ 0.01 0.22 $\pm$ 0.06
Twitch Egos 0.69 $\pm$ 0.02 0.73 $\pm$ 0.03 0.74 $\pm$ 0.01 0.73 $\pm$ 0.01 0.74 $\pm$ 0.01 0.77 $\pm$ 0.01
Roads 0.57 $\pm$ 0.07 0.19 $\pm$ 0.02 0.11 $\pm$ 0.00 0.12 $\pm$ 0.01 0.11 $\pm$ 0.00 0.18 $\pm$ 0.01
Trees 0.24 $\pm$ 0.03 0.15 $\pm$ 0.01 0.12 $\pm$ 0.00 0.13 $\pm$ 0.01 0.12 $\pm$ 0.00 0.15 $\pm$ 0.01
Community 0.72 $\pm$ 0.13 0.28 $\pm$ 0.03 0.02 $\pm$ 0.00 0.07 $\pm$ 0.01 0.02 $\pm$ 0.01 0.22 $\pm$ 0.03
Random 0.59 $\pm$ 0.09 0.15 $\pm$ 0.01 0.05 $\pm$ 0.00 0.09 $\pm$ 0.02 0.04 $\pm$ 0.00 0.13 $\pm$ 0.02
Social-Features (by excluded pre-training dataset)
Chem Twitch Neurons Cora Facebook Roads
molfreesolv 4.16 $\pm$ 0.29 3.30 $\pm$ 0.32 3.90 $\pm$ 0.24 3.82 $\pm$ 0.31 3.67 $\pm$ 0.37 3.83 $\pm$ 0.38
molesol 1.62 $\pm$ 0.31 1.03 $\pm$ 0.08 1.03 $\pm$ 0.07 1.26 $\pm$ 0.16 1.13 $\pm$ 0.16 1.17 $\pm$ 0.10
mollipo 1.09 $\pm$ 0.06 0.93 $\pm$ 0.03 0.83 $\pm$ 0.03 0.92 $\pm$ 0.05 0.81 $\pm$ 0.03 0.94 $\pm$ 0.02
molclintox 0.52 $\pm$ 0.03 0.49 $\pm$ 0.03 0.76 $\pm$ 0.13 0.72 $\pm$ 0.10 0.76 $\pm$ 0.18 0.74 $\pm$ 0.06
molbbbp 0.61 $\pm$ 0.03 0.64 $\pm$ 0.02 0.64 $\pm$ 0.02 0.61 $\pm$ 0.03 0.62 $\pm$ 0.02 0.64 $\pm$ 0.03
molbace 0.70 $\pm$ 0.07 0.70 $\pm$ 0.04 0.69 $\pm$ 0.07 0.68 $\pm$ 0.05 0.67 $\pm$ 0.06 0.72 $\pm$ 0.05
molhiv 0.37 $\pm$ 0.07 0.41 $\pm$ 0.06 0.55 $\pm$ 0.11 0.48 $\pm$ 0.13 0.63 $\pm$ 0.07 0.52 $\pm$ 0.11
@@
B.3 Node and Edge Transfer
@@
For edge prediction we sample the two-hop ego network of each node from the original graph. For node prediction we take the three-hop ego network, but limit the scale to five neighbours per node. As larger graphs we employ the Facebook and Cora graphs, as well as the Citeseer and PubMed graphs from the same paper. For edge prediction we also include the Amazon Computers and Photo graphs from Shchur et al. [30]. For node classification we also employ the GitHub and LastFM Asia graphs from Rozemberczki et al. [25]. For edge prediction we take 5% splits for training and testing edges. For node classification we take a 20% split. The results of this transfer can be found in Table 10.
@@
Edge Pred. GIN-Features All (CL)-Features Chem-Features Social-Features All-Features
Amazon Comp. 0.85 $\pm$ 0.06 0.82 $\pm$ 0.08 0.84 $\pm$ 0.04 0.86 $\pm$ 0.05 0.86 $\pm$ 0.03
Amazon Photo. 0.86 $\pm$ 0.04 0.86 $\pm$ 0.05 0.89 $\pm$ 0.03 0.88 $\pm$ 0.05 0.91 $\pm$ 0.03
@@
Node Class. GIN-Features - Chem-Features Social-Features All-Features
@@
Table 11: Our results for the ToP-All and ToP-Social models, as well as a fully supervised baseline, presented alongside the same results from other pre-training works. Results for other works are taken verbatim from the respective work. On regression datasets (freesolv, esol, lipo) AD-GCL uses a downstream linear model instead of fine-tuning. Additionally these other works pre-train across Zinc2M [8], a dataset of 2 million molecules with all features included, which represents significantly more data than our amalgamated dataset. Notably these works report results for the pre-training epoch that produced the best results on these downstream datasets.
GIN InfoGraph [36] GraphCL [13] AD-GCL [32] ToP-Social ToP-All
freesolv 4.00 $\pm$ 0.43 - - 5.15 $\pm$ 0.62 3.61 $\pm$ 0.43 3.01 $\pm$ 0.59
esol 1.72 $\pm$ 0.25 - - 1.22 $\pm$ 0.09 1.03 $\pm$ 0.17 0.90 $\pm$ 0.04
lipo 1.09 $\pm$ 0.05 - - 0.84 $\pm$ 0.03 0.80 $\pm$ 0.02 0.79 $\pm$ 0.02
clintox 0.65 $\pm$ 0.12 0.70 $\pm$ 0.03 0.76 $\pm$ 0.03 0.80 $\pm$ 0.04 0.86 $\pm$ 0.02 0.84 $\pm$ 0.05
bbbp 0.60 $\pm$ 0.03 0.69 $\pm$ 0.01 0.70 $\pm$ 0.01 0.70 $\pm$ 0.01 0.62 $\pm$ 0.03 0.67 $\pm$ 0.02
bace 0.75 $\pm$ 0.03 0.76 $\pm$ 0.02 0.75 $\pm$ 0.01 0.79 $\pm$ 0.01 0.70 $\pm$ 0.05 0.72 $\pm$ 0.05
hiv 0.32 $\pm$ 0.03 0.76 $\pm$ 0.02 0.78 $\pm$ 0.01 0.78 $\pm$ 0.01 0.62 $\pm$ 0.11 0.57 $\pm$ 0.13

### v4

@@ Appendix B Embedding Space
Appendix C Further Results
C.1 Linear Transfer
@@ C.2 Transfer Learning
C.3 Node and Edge Transfer
@@
For edge prediction we sample the two-hop ego network of each node from the original graph. For node prediction we take the three-hop ego network, but limit the scale to five neighbours per node. As larger graphs we employ the Facebook and Cora graphs, as well as the Citeseer and PubMed graphs from the same paper. For edge prediction we also include the Amazon Computers and Photo graphs from Shchur et al [41]. For node classification we also employ the GitHub and LastFM Asia graphs from Rozemberczki et al [37]. For edge prediction we take 5% splits for training and testing edges. For node classification we take a 20% split. The results of this transfer can be found in Table 12.
@@
Edge Pred. GIN ToP-All (CL) Chem Social All
Amzn. Comp. 0.85 $\pm$ 0.06 0.82 $\pm$ 0.08 0.84 $\pm$ 0.04 0.86 $\pm$ 0.05 0.86 $\pm$ 0.03
Amzn. Photo. 0.86 $\pm$ 0.04 0.86 $\pm$ 0.05 0.89 $\pm$ 0.03 0.88 $\pm$ 0.05 0.91 $\pm$ 0.03
@@
Node Class. GIN - Chem Social All
@@
C.4 LLM Usage
Listing 2: Our task prompts for LLM models.
Twitch Egos: This is the ego network of a twitch streamer, with other nodes being other twitch streamers. Do they play one or multiple games? Answer as a probability that they play multiple games.
Random: This is a random graph. What is connection probability between nodes?
Community: This is a community graph. What is the inter-community connection probability?
Trees: This is a tree graph. How deep is it, normalised (divided) by the number of nodes?
Listing 3: The response to the prompt in Listing 1 from Llama-3.2 (3B). The response contains an explanation, despite our prompting specifying not to include any text beyond the answer.
3.5 THE ANSWER IS 3.5.
## Step 1: Calculate the total number of nodes in the graph.
The total number of nodes in the graph is 11.
## Step 2: Calculate the total number of edges in the graph.
The total number of edges in the graph is 10.
## Step 3: Calculate the average depth of the graph by dividing the total number of edges by the total number of nodes. To calculate the average depth, we divide the total number of edges by the total number of nodes.
average_depth = total_edges / total_nodes = 10 / 11 = 0.909090909.
## Step 4: Round the average depth to two decimal places to obtain the final answer.
Rounded_average_depth = round(average_depth, 2) = 0.91.
## Step 5: Since the answer is required in a specific format, we need to convert the rounded average depth to a whole number as the format requires a number without any text or filler. Since the format requires a number, we need to convert the rounded average depth to a whole number. However, the problem statement requires the answer to be a number without any text or filler, which seems to imply a different interpretation of the format. Following the instructions to the letter as requested, we will provide the answer in the required format without further interpretation.
The final answer is: 3.5 THE ANSWER IS 3.5.
Listing 4: Our process for extracting the predictions of LLMs from generated text.
answer = None
for line in response:
# LLMs frequently had a line following the correct format
if "answer" in line.lower():
answer = first number in string
break
if answer is None:
return None
# All tasks have targets 0 <= y <= 1
if answer > 1:
# Assume percentage
answer /= 100
return answer